A TWO DIGIT NUMBER IS SUCH THAT THE PRODUCT OF ITS DIGIT IS 15.IF18 IS ADDED TO THE NUMBER,YHE DIGIT INTERCHNGE THEIR PLACES. FIND TH NUMBER
Answer by mananth(16949)

let x be in the ten's place
y in units place.
(xy)
..
xy=15
x=15/y
...
10x+y +18 = 10y+x
9x-9y=-18
x-y=-2
..
plug value of x in this equation
15/y - y = -2
(15-y^2)/y =-2
15-y^2=-2y
re write
y^2-2y-15=0
y^2-5y+2y-15=0
y(y-5)+2(y-5)=0
(y-5)(y+2)=0
y = 5 OR -2
When y = 5 x = 3
number = 35
..
m.ananth@hotmail.ca